OpenStudy (anonymous):

10 out 30 is what %?

myininaya (myininaya):

\[\frac{10}{30}=\frac{1}{3}=\frac{1}{3} \cdot 1=\frac{1}{3} \cdot 100\%=\frac{100}{3} \%\]

OpenStudy (kinggeorge):

\[{10 \over 30} = {x \over 100}\] so \[{30 x = 1,000}\] so \[x = {1000 \over 30} = {100 \over 3}\] or x = 33.33%